theyadoreshayyy:

Find the x‐value for point C such that AC and BC form a 2:3 ratio.

theyadoreshayyy:

theyadoreshayyy:

@Tyrion

theyadoreshayyy:

@Conner4789

jakfishman:

@snowflake0531

minustempo:

It's a fact that if AC:BC = 2/3, then x-diff of A to C : x-diff of B to C = 2:3, and same with the y value

minustempo:

though the problem wants the x value of C

minustempo:

So the x-value of A is -3, and the x-value of B is 3 let the x-value of C just be x

minustempo:

So the distance from -3 to x : the distance from x to 3 = 2/3

minustempo:

so here, we get (x - (-3)) : (3 - x) = 2:3

minustempo:

So (x+3) : (3-x) = 2:3 which you can rewrite as fractions \[(x+3)/(3-x) = 2/3\]

minustempo:

Cross multiplying gets you 3(x+3) = 2(3-x) 3x+9 = 6-2x 5x = -3 x = -0.6

minustempo:

hope that helped!

theyadoreshayyy:

i did thank you soooo much